# Geometric Embeddability of Complexes is $\exists\mathbb{R}$-complete

### 05 Aug 2021 v1

@@ Abstract
We show that the decision problem of determining whether a given (abstract simplicial) $k$-complex has a geometric embedding in $\mathbb{R}^{d}$ is complete for the Existential Theory of the Reals for all $d\geq 3$ and $k\in\{d-1,d\}$. This implies that the problem is polynomial time equivalent to determining whether a polynomial equation system has a real root. Moreover, this implies NP-hardness and constitutes the first hardness results for the algorithmic problem of geometric embedding (abstract simplicial) complexes.
Figure 1: Illustration of different embeddings of complexes; figure taken from Bing [8, Annals of Mathematics 19591959].
@@ 1 Introduction
Since the dawn of the last century, much attention has been devoted to studying embeddings of complexes [29, 8, 20, 27, 28, 37, 48, 59]. Typical types of embeddings include geometric (also referred to as linear), piecewise linear (PL), and topological embeddings, see also Figure 1. For formal definitions, we refer to Section 1.2; here we give an illustrative example. Embeddings of a $1$-complex in the plane correspond to drawings of a graph in the plane. In a topological embedding, each edge is represented by a Jordan arc, in a PL embedding it is a concatenation of a finite number of segments, and in a geometric embedding each edge is a segment. Unlike in higher dimensions, for the embeddability of complexes in the plane all three notions coincide.
@@
We are interested in the problem of deciding whether a given $k$-complex has a linear/piecewise linear/topological embedding in $\mathbb{R}^{d}$. Several necessary and sufficient conditions are easy to identify and have been known for many decades. For instance, a $k$-simplex requires $(k+1)$ points in general position in $\mathbb{R}^{d}$ and, thus, $k\leq d$ is an obvious necessary condition. Moreover, it is straight-forward to verify that every set of $n$ points in $\mathbb{R}^{3}$ in general position allows for a geometric embedding of any 1-complex on $n$ vertices, i.e., straight-line drawing of a (complete) graph. Indeed, this fact generalizes to higher dimensions: every $k$-complex embeds (even linearly) in $\mathbb{R}^{2k+1}$ [37]. Van Kampen and Flores [24, 51, 59] showed that this bound is tight by providing $k$-complexes that do not topologically embed into $\mathbb{R}^{2k}$. For some time, it was believed that the existence of a topological embedding also implies the existence of a geometric embedding, e.g., Grünbaum conjectured that if a $k$-dimensional simplicial complex is topologically embeddable in $\mathbb{R}^{2k}$ then it also geometrically embeds in $\mathbb{R}^{2k}$ [27]. However, this was consequently disproven. In particular, for every $k\geq 2$ and $k+1\leq d\leq 2k$, there exist $k$-complexes that have a PL embedding, but no geometric embedding in $\mathbb{R}^{d}$ [9, 10]. In contrast, PL and topological embeddability coincides in many cases, e.g., if $d\leq 3$ [8, 43] or $d-k\geq 3$ [11]. There are many further necessary and sufficient conditions for geometric embeddings [6, 41, 42, 51, 56, 57] and PL or/and topological embeddings [19, 25, 44, 48, 58].
@@
In recent years, the algorithmic complexity of deciding whether or not a given complex is embeddable gained attention. In the absence of a complete characterization, an efficient algorithm is the best tool to decide embeddability. For instance, embeddings of 1-complexes in the plane corresponds to testing graph planarity which is known to be polynomial time decidable [30]. Conversely, the proven non-existence of efficient algorithms may offer a rigorous proof that a complete characterization is impossible. To give a concrete example, let $\textsc{Embed}_{k\rightarrow d}$ denote the algorithmic problem of determining whether a given $k$-complex has a PL embedding in $\mathbb{R}^{d}$. Because $\textsc{Embed}_{4\rightarrow 5}$ is known to be undecidable [23, 53], we have a proof that there does not exist an efficient algorithm for $\textsc{Embed}_{4\rightarrow 5}$ – even without any complexity assumptions (such as $\mathrm{NP}\neq\mathrm{P}$ or similar).
@@
More recently, there have been several breakthroughs concerning the PL-embeddability. For an overview of the state of the art, consider Table 1.
Table 1: Overview for PL
@@
In dimensions $d\geq 4$, the decision problem $\textsc{Embed}_{k\rightarrow d}$ is polynomial-time decidable for $k<\nicefrac{{2}}{{3}}(d-1)$ [15, 12, 14, 31] and NP-hard for all remaining cases [35], i.e., for all $k\geq\nicefrac{{2}}{{3}}(d-1)$. For $d\leq 5$ and $k\in\{d,d-1\}$, $\textsc{Embed}_{k\rightarrow d}$ is even known to be undecidable; for all other NP-hard cases and $d\geq 4$ decidability is unknown. For the case $d=3$, Matoušek, Sedgwick, Tancer, and Wagner have shown that $\textsc{Embed}_{2\rightarrow 3}$ and $\textsc{Embed}_{3\rightarrow 3}$ are decidable [34] and de Mesmay, Rieck, Sedgwick, and Tancer have proved NP-hardness [38].
@@
Building upon [35], Skopenkov and Tancer [55] proved NP-hardness for a relaxed notion called almost (PL/topological) embeddability where disjoint hyperedges must be mapped to disjoint objects. More precisely, they showed that recognizing almost embeddability of $k$-complexes in $\mathbb{R}^{d}$ is $\mathrm{NP}$-hard for all $d,k\geq 2$ such that $d\pmod{3}=1$ and $\nicefrac{{2}}{{3}}(d-1)\leq k\leq d$.
@@
The analogous questions for geometric embeddings are wide open. Let $\textsc{GEM}_{k\rightarrow d}$ denote the algorithmic problem of determining whether a given $k$-complex has a geometric embedding in $\mathbb{R}^{d}$. In contrast to PL-embeddability, however, it is easy to see that $\textsc{GEM}_{k\rightarrow d}$} is decidable for all $k,d$. In analogy to the PL-embeddings, $\mathrm{NP}$-hardness has been conjectured by Skopenkov.
Conjecture ([54], Conjecture 3.2.2).
@@
$\max\{3,k\}\leq d\leq\nicefrac{{3}}{{2}}k+1.$
Note that these pairs correspond to the NP-hard cases of $\textsc{Embed}_{k\rightarrow d}$.
Results.
In this work, we present the first results for any non-trivial entry with $d\geq 3$. More precisely, we establish the exact computational complexity of $\textsc{GEM}_{k\rightarrow d}$ for all values $d\geq 3$ and $k\in\{d,d-1\}$, hereby confirming the conjecture for these cases. This includes a complete understanding of the most intriguing entries with $d=3$. Table 2 summarizes the current knowledge on the computational complexity of $\textsc{GEM}_{k\rightarrow d}$.
@@ Results.
For every $d\geq 3$ and every $k\in\{d,d-1\}$, the decision problem $\textsc{GEM}_{k\rightarrow d}$ is $\exists\mathbb{R}$-complete.
Moreover, the statement remains true even if a PL-embedding is given.
@@
As $\mathrm{NP}\subseteq\exists\mathbb{R}$, our result confirms the conjecture by Skopenkov for the corresponding values of $k$ and $d$. Moreover, if $\mathrm{NP}\neq\exists\mathbb{R}$, $\textsc{GEM}_{k\rightarrow d}$ cannot be tackled with well developed tools for $\mathrm{NP}$-complete problems such as SAT solvers and ILPs. For more details, we refer to Section 1.1.
As a matter of fact, our proof implies that distinguishing between $k$-complexes with PL and geometric embeddings in $\mathbb{R}^{d}$ is complete for $\exists\mathbb{R}$.
Geometric embeddings of complexes can also be viewed as a representation of hypergraphs in which every edge is represented by a simplex. Of particular interest is the case of uniform hypergraphs where all hyperedges have the same number of elements. Thus, in the language of hypergraphs, our result reads as follows.
@@
For all $d\geq 3$ and every $k\in\{d,d-1\}$, deciding whether a $(k+1)$-uniform hypergraph has a simplicial representation in $\mathbb{R}^{d}$ is $\exists\mathbb{R}$-complete.
@@
The main idea to prove hardness of $\textsc{GEM}_{2\rightarrow 3}$} is to reduce from the Stretchability problem. In Stretchability, we are given a set of pseudolines (curves) in the plane and we are asked to decide if there are straight lines that intersect each other in the same combinatorial pattern as the pseudolines. See below for a formal definition. Given a pseudoline arrangement $L$, we construct a $2$-complex $C$, such that $C$ is realizable in $\mathbb{R}^{3}$ if and only if $L$ is stretchable.
In our construction each pseudoline is represented by a special edge. We wrap tunnels around the special edges and all tunnels are connected by triangles to a special apex vertex $u$. It will be straightforward to verify that if $L$ is stretchable, then the resulting simplex $C$ is embeddable. The other direction is more subtle: We show that if $C$ is embeddable, then seen from $u$, the special edges will appear as a realization of $L$ using straight lines, and hence $L$ is stretchable. The idea behind the proof is that the tunnels and the apex $u$ force together the correct combinatorics of the intersection pattern of the special edges seen from $u$.
@@
In order to show hardness of $\textsc{GEM}_{3\rightarrow 3}$}, we use a similar construction, but where we “fatten” each triangle to a tetrahedron, by adding extra vertices.
@@
We finally make a dimension reduction by proving that $\textsc{GEM}_{k\rightarrow d}$} reduces to $\textsc{GEM}_{k+1\rightarrow d+1}$}. Given a $k$-complex $C$, we create a $(k+1)$-complex $C^{+}$ in the following way. The vertices of $C^{+}$ are those of $C$ and two extra vertices $a$ and $b$. For each edge $e$ of $C$, we have the additional edges $e\cup\{a\}$ and $e\cup\{b\}$ in $C^{+}$. We prove that $C$ geometrically embeds in $\mathbb{R}^{d}$ if and only if $C^{+}$ geometrically embeds in $\mathbb{R}^{d+1}$.
@@
Thus, we show that deciding whether a given complex with a PL-embedding also allows a geometric embedding is $\exists\mathbb{R}$-complete, i.e., distinguishing PL-embeddable and geometric embeddable complexes is $\exists\mathbb{R}$-complete.
@@ 1.1 Existential Theory of the Reals
The first inclusion follows from the definition of $\exists\mathbb{R}$. Showing the second inclusion was first established by Canny in his seminal paper [16]. The complexity class $\exists\mathbb{R}$ gains its significance because a number of well-studied problems from different areas of theoretical computer science have been shown to be complete for this class.
Famous examples from discrete geometry are the recognition of geometric structures, such as unit disk graphs [36], segment intersection graphs [33], stretchability [40, 50], and order type realizability [33]. Other $\exists\mathbb{R}$-complete problems are related to graph drawing [32], Nash-Equilibria [7, 26], geometric packing [5], the art gallery problem [3], non-negative matrix factorization [49], polytopes [21, 46], geometric linkage constructions [1], training neural networks [4], visibility graphs [18], continuous CSPs [39], and convex covers [2]. The fascination for the complexity class stems not merely from the number of $\exists\mathbb{R}$-complete problems but from the large scope of seemingly unrelated $\exists\mathbb{R}$-complete problems. We refer the reader to the lecture notes by Matoušek [33] and surveys by Schaefer [47] and Cardinal [17] for more information on the complexity class $\exists\mathbb{R}$.

### 05 Nov 2021 v2

@@ Abstract
We show that the decision problem of determining whether a given (abstract simplicial) $k$-complex has a geometric embedding in $\mathbb{R}^{d}$ is complete for the Existential Theory of the Reals for all $d\geq 3$ and $k\in\{d-1,d\}$. This implies that the problem is polynomial time equivalent to determining whether a polynomial equation system has a real solution. Moreover, this implies NP-hardness and constitutes the first hardness result for the algorithmic problem of geometric embedding (abstract simplicial) complexes.
Figure 1: Illustration of different embeddings of a complex; figure taken from Bing [undefg, Annals of Mathematics 19591959].
@@ 1 Introduction
Since the dawn of the last century, much attention has been devoted to studying embeddings of complexes [undefac, undefg, undeft, undefaa, undefab, undefal, undefay, undefaak]. Typical types of embeddings include geometric (also referred to as linear), piecewise linear (PL), and topological embeddings, see also Figure 1. For formal definitions, we refer to Section 1.2; here we give an illustrative example. Embeddings of a $1$-complex in the plane correspond to drawings of a graph in the plane. In a topological embedding, each edge is represented by a Jordan arc, in a PL embedding it is a concatenation of a finite number of segments, and in a geometric embedding each edge is represented by a segment. Unlike in higher dimensions, for the embeddability of complexes in the plane all three notions coincide.
@@
We are interested in the problem of deciding whether a given $k$-complex has a linear/piecewise linear/topological embedding in $\mathbb{R}^{d}$. Several necessary and sufficient conditions are easy to identify and have been known for many decades. For instance, a $k$-simplex requires $k+1$ points in general position in $\mathbb{R}^{d}$ and, thus, $k\leq d$ is an obvious necessary condition. Moreover, it is straight-forward to verify that every set of $n$ points in $\mathbb{R}^{3}$ in general position allows for a geometric embedding of any 1-complex on $n$ vertices, i.e., the points are the vertices of a straight-line drawing of a (complete) graph. Indeed, this fact generalizes to higher dimensions: every $k$-complex embeds (even linearly) in $\mathbb{R}^{2k+1}$ [undefal]. Van Kampen and Flores [undefx, undefaae, undefaak] showed that this bound is tight by providing $k$-complexes that do not topologically embed into $\mathbb{R}^{2k}$. For some time, it was believed that the existence of a topological embedding also implies the existence of a geometric embedding, e.g., Grünbaum conjectured that if a $k$-complex topologically embeds in $\mathbb{R}^{2k}$, then it also geometrically embeds in $\mathbb{R}^{2k}$ [undefaa]. However, this was later disproven. In particular, for every $k,d\geq 2$ with $k+1\leq d\leq 2k$, there exist $k$-complexes that have a PL embedding in $\mathbb{R}^{d}$, but no geometric embedding in $\mathbb{R}^{d}$ [undefj, undefh, undefi]. In contrast, PL and topological embeddability coincides in many cases, e.g., if $d\leq 3$ [undefg, undefar] or $d-k\geq 3$ [undefk]. There are many further necessary and sufficient conditions known for geometric embeddings [undefe, undefap, undefaq, undefaae, undefaah, undefaai] and PL or/and topological embeddings [undefs, undefy, undefas, undefay, undefaaj, undefaag].
@@
In recent years, the algorithmic complexity of deciding whether or not a given complex is embeddable gained attention. In the absence of a complete characterization, an efficient algorithm is the best tool to decide embeddability. For instance, deciding whether a 1-complex embeds in the plane corresponds to testing graph planarity and is thus polynomial time decidable [undefad]. Conversely, the proven non-existence of efficient algorithms may offer a rigorous proof that a complete characterization is impossible. To give a concrete example, let $\textsc{Embed}_{k\rightarrow d}$ denote the algorithmic problem of determining whether a given $k$-complex has a PL embedding in $\mathbb{R}^{d}$. Because $\textsc{Embed}_{4\rightarrow 5}$ is known to be undecidable [undefw, undefaac], we have a proof that there does not exist an efficient algorithm for $\textsc{Embed}_{4\rightarrow 5}$ – even without any complexity assumptions (such as $\mathrm{NP}\neq\mathrm{P}$ or similar).
@@
More recently, there have been several breakthroughs concerning the PL embeddability. For an overview of the state of the art, consider Table 1.
Table 1: Overview of the complexity of $\textsc{Embed}_{k\rightarrow d}$.
@@
In dimensions $d\geq 4$, the decision problem $\textsc{Embed}_{k\rightarrow d}$ is polynomial-time decidable for $k<\nicefrac{{2}}{{3}}\cdot(d-1)$ [undefo, undefl, undefn, undefaf] and NP-hard for all remaining cases [undefaj], i.e., for all $k\geq\nicefrac{{2}}{{3}}\cdot(d-1)$. For $d\geq 5$ and $k\in\{d-1,d\}$, $\textsc{Embed}_{k\rightarrow d}$ is even known to be undecidable; for all other NP-hard cases and $d\geq 4$ decidability is unknown. For the case $d=3$, Matoušek, Sedgwick, Tancer, and Wagner have shown that $\textsc{Embed}_{2\rightarrow 3}$ and $\textsc{Embed}_{3\rightarrow 3}$ are decidable [undefai] and de Mesmay, Rieck, Sedgwick, and Tancer have proved NP-hardness [undefam].
@@
Building upon [undefaj], Skopenkov and Tancer [undefaaf] proved NP-hardness for a relaxed notion called almost (PL/topological) embeddability where it is only required that disjoint sets must be mapped to disjoint objects, i.e., this notions allows that two edges incident to a common vertex cross in an interior point. More precisely, they showed that recognizing almost embeddability of $k$-complexes in $\mathbb{R}^{d}$ is $\mathrm{NP}$-hard for all $d,k\geq 2$ such that $d\pmod{3}=1$ and $\nicefrac{{2}}{{3}}\cdot(d-1)\leq k\leq d$.
@@
The analogous questions for geometric embeddings are wide open. Let $\textsc{GEM}_{k\rightarrow d}$ denote the algorithmic problem of determining whether a given $k$-complex has a geometric embedding in $\mathbb{R}^{d}$. In contrast to PL embeddability, however, it is easy to see that $\textsc{GEM}_{k\rightarrow d}$} is decidable for all $k,d$, since every instance can be expressed as a sentence in the first order theory of the reals, which is decidable. In analogy to the PL embeddings, $\mathrm{NP}$-hardness has been conjectured by Skopenkov.
Conjecture ([undefaad], Conjecture 3.2.2).
@@
$\max\{3,k\}\leq d\leq\nicefrac{{3}}{{2}}\cdot k+1.$
Note that these parameters correspond to the NP-hard cases of $\textsc{Embed}_{k\rightarrow d}$, see also Table 1. Cardinal [undefq, Section 4] mentions $\textsc{GEM}_{2\rightarrow 3}$} as an interesting open problem. The closely related question of polyhedral complexes, posed in the Handbook of Discrete and Computational Geometry, reads as follows: When is a given finite poset isomorphic to the face poset of some polyhedral complex in a given space $\mathbb{R}^{d}$[undefax, Problem 20.1.1]. Note that simplicial complexes are special cases of polyhedral complexes, because each simplex is a basic polyhedron. The recognition of polyhedral complexes (with triangles and quadrangles) in $\mathbb{R}^{3}$ has been claimed to be $\exists\mathbb{R}$-complete [undefq, Theorem 5]. Focussing on convex polytopes, Richter-Gebert proved that recognizing convex polytopes in $\mathbb{R}^{4}$ is $\exists\mathbb{R}$-complete [undefau, undefav].
Our Results.
In this work, we present the first results concerning Skopenkov’s conjecture for any non-trivial entry with $d\geq 3$. More precisely, we establish the exact computational complexity of $\textsc{GEM}_{k\rightarrow d}$ for all values $d\geq 3$ and $k\in\{d-1,d\}$, hereby confirming the conjecture for these cases. This includes a complete understanding of the most intriguing entries with $d=3$. Note that this also answers the computational aspects of the question from the Handbook of Discrete and Computational Geometry. Table 2 summarizes the current knowledge on the computational complexity of $\textsc{GEM}_{k\rightarrow d}$.
@@
For every $d\geq 3$ and each $k\in\{d-1,d\}$, the decision problem $\textsc{GEM}_{k\rightarrow d}$ is $\exists\mathbb{R}$-complete. Moreover, the statement remains true even if a PL embedding is given.
@@
Our proof implies that distinguishing between $k$-complexes with PL and geometric embeddings in $\mathbb{R}^{d}$ is complete for $\exists\mathbb{R}$. Because $\mathrm{NP}\subseteq\exists\mathbb{R}$, our result confirms the conjecture by Skopenkov for the corresponding values of $k$ and $d$. Moreover, if $\mathrm{NP}\neq\exists\mathbb{R}$, the problem $\textsc{GEM}_{k\rightarrow d}$ cannot be tackled with well developed tools for $\mathrm{NP}$-complete problems such as SAT and ILP solvers. For more details, we refer to Section 1.1.
A geometric embedding of a complex can also be viewed as a simplicial representation of a hypergraph, i.e., a representation of a hypergraph in which every hyperedge is represented by a simplex. Of particular interest is the case of uniform hypergraphs where all hyperedges have the same number of elements. Thus, in the language of hypergraphs, our result reads as follows.
@@
For all $d\geq 3$ and every $k\in\{d-1,d\}$, deciding whether a $(k+1)$-uniform hypergraph has a simplicial representation in $\mathbb{R}^{d}$ is $\exists\mathbb{R}$-complete.
@@
Our proof of Theorem 1 consists of three steps: Establishing $\exists\mathbb{R}$-membership, showing $\exists\mathbb{R}$-hardness in $\mathbb{R}^{3}$, i.e., of $\textsc{GEM}_{2\rightarrow 3}$ and $\textsc{GEM}_{3\rightarrow 3}$, and reducing $\textsc{GEM}_{k\rightarrow d}$ to $\textsc{GEM}_{k+1\rightarrow d+1}$. The core of the proof lies in establishing hardness of $\textsc{GEM}_{2\rightarrow 3}$}.
@@
The main idea to prove hardness of $\textsc{GEM}_{2\rightarrow 3}$} is to reduce from the problem Stretchability. In Stretchability, we are given an arrangement of pseudolines (curves) in the plane and we are asked to decide whether there exists a set of straight lines that has the same combinatorial pattern as the pseudoline arrangement, see Figure 2(a) for an illustration and Section 1.2 for a formal definition. Given a pseudoline arrangement $L$, we construct a $2$-complex $C$ which has a geometric embedding in $\mathbb{R}^{3}$ if and only if $L$ is stretchable.
Figure 2: (a) We start with a pseudoline arrangement $L$. (b) We add a triangle containing all intersections of $L$. (c) Each pseudoline is represented by a special edge that is surrounded by a tunnel. (d) Each tunnel consists of tunnel sections. (e) For the crossings of the special edges, we identify parts of the tunnels. (f) We add an apex $u$ and insert triangles to the visible parts of the construction; we enhance the neighborhood of the apex to an essentially 3-connnected graph depicted in (i). Together, (e), (f) and (i) are crucial to enforce the correct combinatorics. (g) In the correctness proof, we use a small sphere around the apex and the projection of each special edge onto the sphere. (h) We will argue that the combinatorics on the sphere are correct and then project the special edges onto a plane. This will give a stretched arrangement. (i) The neighborhood graph of the apex $u$.
On a high level, our construction of $C$ goes along the following lines: We add a helper triangle that contains all intersections of the pseudolines, see Figure 2(b). We place each pseudoline in $\mathbb{R}^{3}$ and replace it by a special edge of the complex $C$. We surround the special edges by tunnels, see Figure 2(c) and (d). For each crossing in $L$, we glue the corresponding tunnel sections together, see Figure 2(e). At last, we insert an apex $u$ high above that is connected to all visible tunnel parts, see Figure 2(f) and we insert additional objects in order to ensure that the neighborhood of $u$ is an essentially 3-connected graph, Figure 2(i).
It is relatively straightforward to verify that if $L$ is stretchable, then the complex $C$ embeds geometrically into $\mathbb{R}^{3}$. The other direction requires more care and work: We show that a geometric embedding of $C$ induces a line arrangement with the same combinatorics as $L$. The idea of the proof is to consider a small sphere around the apex $u$ and to project its neighborhood and the special edges onto the sphere, see Figure 2(g). Because the neighborhood graph of $u$ is essentially 3-connected by construction, all its crossing-free drawings on the sphere are equivalent. This is an crucial property to show that each special edge lies in the projection of its tunnel roof (when restriction the attention to an interesting part within the helper triangle). We remark, that our proof does not show this explicitly. Instead, we establish some even stronger properties. As a consequence, the projection of the tunnels have the intended combinatorics and thus also the special edges. At last, we project the arcs from the sphere onto a plane, see Figure 2(h). In this way, we obtain a line arrangement with the same combinatorics as $L$.
@@
In order to show hardness of $\textsc{GEM}_{3\rightarrow 3}$}, we use a similar construction, in which we “fatten” each triangle to a tetrahedron, by adding extra vertices.
@@
We finally present a dimension reduction, i.e., we reduce $\textsc{GEM}_{k\rightarrow d}$} to $\textsc{GEM}_{k+1\rightarrow d+1}$}. Given a $k$-complex $C$, we create a $(k+1)$-complex $C^{+}$ that contains $C$ and has two additional vertices $a$ and $b$. Moreover, for each subset $e$ of $C$, $C^{+}$ has the additional subsets $e\cup\{a\}$ and $e\cup\{b\}$. We prove that $C$ geometrically embeds in $\mathbb{R}^{d}$ if and only if $C^{+}$ geometrically embeds in $\mathbb{R}^{d+1}$.
@@
In this way, we show that distinguishing PL embeddable and geometrically embeddable complexes is $\exists\mathbb{R}$-complete.
@@ 1.1 Existential Theory of the Reals
The first inclusion follows from the definition of $\exists\mathbb{R}$. Showing the second inclusion was first established by Canny in his seminal paper [undefp]. The complexity class $\exists\mathbb{R}$ gains its significance because a number of well-studied problems from different areas of theoretical computer science have been shown to be complete for this class.
Famous examples from discrete geometry are the recognition of geometric structures, such as unit disk graphs [undefak], segment intersection graphs [undefah], Stretchability [undefao, undefaaa], and order type realizability [undefah]. Other $\exists\mathbb{R}$-complete problems are related to graph drawing [undefag], Nash-Equilibria [undeff, undefz], geometric packing [undefd], the art gallery problem [undefb], non-negative matrix factorization [undefaz], polytopes [undefu, undefav], geometric linkage constructions [undef], training neural networks [undefc], visibility graphs [undefr], continuous constraint satisfaction problems [undefan], and convex covers [undefa]. The fascination for the complexity class stems not merely from the number of $\exists\mathbb{R}$-complete problems but from the large scope of seemingly unrelated $\exists\mathbb{R}$-complete problems. We refer the reader to the lecture notes by Matoušek [undefah] and surveys by Schaefer [undefaw] and Cardinal [undefq] for more information on the complexity class $\exists\mathbb{R}$.